Consider a city that has cell phone case stands operating throughout the midtown area. Suppose each vendor has a marginal cost of $5.00 per case and no fixed cost. Suppose the maximum number of cell phone cases that any one vendor can sell is 70 per day.

  • If the price of a cell phone case is $15.00, how many cases does each vendor want to sell? B. If the industry is perfectly competitive, will the price remain $15.00 per case? If not, what will the price be? Why?
  • If each vendor sells exactly 70 cell phone cases a day and the demand for cell phone cases for vendors in the city is Q=1000-20P, how many vendors are there?
  • The domestic supply and demand curves for washing machines are as follows:

Supply: P= 2800+5Q Demand: P=4300-5Q

where P is the price in dollars and the Q is the quantity in millions. The U.S. is a small producer in the world washing machine market. Where the current price (which will not be affected by anything we do) is $ 3,000.

Congress is considering a tariff of $500.

A. Calculate and graph all points for the domestic market for washing machines price and quantity equilibrium.

B. Find the domestic quantity demanded and supplied of washing machines that will result if the price imposition of $3,000 is imposed. Show on graph. Explain.

  • Find the domestic quantity demanded and supplied of washing machines that will result if the $500 tariff is imposed. Show on graph. Explain.
  • Compute government revenue from the tariff.
  • Illustrate graphically

Suppose that a competitive firms marginal cost of producing output q is given by

MC(q)= 70+6q

Assume that the market price of the firm’s product is $145.

  • At what level of output will the firm produce?
  • How much is the firm’s producer surplus?
  • Illustrate graphically profit maximization point and producer surplus.
  • Illustrate this market at a loss. Explain.
  • Graphically illustrate a perfectly competitive firm and a non-perfectly competitive firm side by side. Explain the differences.
  • Illustrate graphically a monopolistic competitive firm at a above normal, normal and zero economic profit. (Three separate graphs)
Answer #1

marginal cost = $5.00 per case

a) If price of phone case = $15.00, then each vendor would like to sell the maximum it can in a day, that is 70 cases. This is because price that they receive for each case is more than the MC that they have to bear.

B) In perfect competition, Price would become equal to MC = $5 per case instead of $15 per case. This is because at price of $15, there is a profit in selling cases. Seeing these profits, other sellers would enter into the market and hence, supply would increase and the price would fall. This would keep on happening till price becomes equal to MC and there is no longer any incentive for any other seller to enter into the market.

c) Each sellers sells 70 cases a day.

Demand for cases:  Q=1000-20P

It ia given in part a) that P= $15

When market price = $15, then Q = 1000-20*15 = 700 units.

700 units is the total market demand.

If each seller sells 70 cases, then number of sellers = total market demand/ quantity sold by each seller = 700/70 = 10

As per the HOMEWORKLIB RULES, first question has to be answered. Please put the remaining questions as a separate post. Thank you.